# Insights of Transitions to Thermoacoustic Instability in Inverse Diffusion Flame using Multifractal Detrended Fluctuation Analysis

### v1

@@ Abstract
Keywords: Inverse diffusion flame, Thermoacosutic instability, Multifractal, Hurst exponent, Transition in dynamics.
preprint: APS/123-QED
@@ I Introduction
“Rijke tube” [1, 2] discovered by Professor P.L. Rijke in 1859, is a cylindrical tube with open ends that contains a heat source. This unique configuration generates two distinct types of flames, influenced by the introduction of oxidizer and fuel: normalized diffusion flame (NDF) [3] and inverse diffusion flame (IDF) [4, 5]. In the NDF configuration, fuel is injected at the center, surrounded by an oxidizer. In contrast, the IDF configuration has a coaxial setup where the fuel surrounds a centralized flow of oxidizer. While both configurations provide stable operations, IDF is generally considered more suitable for practical applications because it results in lower emissions of thermal $NO_{x}$ and soot [6, 7, 8, 8], is less susceptible to lean blowout [9], and offers a wider stability margin [10]. In IDF configuration, the central air jet stabilizes the flame and enhances flame stability by creating negative pressure along the flame axis.
@@
An IDF consists of two distinct zones: the flame torch and the flame base. These zones are connected by a section called the flame neck. In the flame neck region, intense mixing occurs due to the controlled velocity of the air jet, allowing air and unburned fuel to mix effectively before entering the flame torch zone [11]. Furthermore, an IDF combines the benefits of both normal diffusion flames and premixed flames. Therefore, the IDF concept can offer an alternative approach to both lean premixed combustion and normal diffusion flames in applications such as rocket engine combustion, staged combustion, industrial burners, and aviation [12]. IDF is also beneficial in heating applications where a single flame needs to have a higher thermal load (i.e., a higher heat transfer rate) [12]. Compared to typical diffusion flames or premixed flames, IDF is cleaner and safer for these types of applications.
@@
However, the IDF configuration converts heat into sound by promoting self-sustaining and self-amplifying standing waves. This phenomenon is highly relevant to the safety of combustion chambers in industrial burners [13], rocket propulsion  [14, 15, 16], etc. The presence of high-amplitude acoustic waves from thermoacoustic instability can damage the structural integrity of engines [13]. This instability often manifests as loud noise, known as combustion roar, which is undesirable in both industrial and consumer applications and can lead to mission failures.  [17]. Additionally, thermoacoustic isntability can disrupt energy conversion efficiency, resulting in higher fuel consumption and lower output power. In some cases, it may also increase pollutant emissions, particularly in industrial combustion systems. Thus, the challenges of thermoacoustic instability in gas turbine engines, rocket motors, industrial boilers, oxygen-fuel Combustion and industrial process control remain significant for researchers in the twenty-first century [18, 19, 20]. Consequently, investigating thermoacoustic instability across various engine combustors, power generation systems, and aerospace applications has become a key area of focus for scientists. In this study, we examine the critical issues of thermoacoustic instability in an IDF combustor.
@@
Earlier, a significant amount of research had been conducted on IDF to understand their characteristics. Recently, Sun et al. [21] employed the Reynolds-averaged Navier-Stokes (RANS) numerical method to analyze the combustion performance of an IDF combined with air swirling. Their results demonstrated that the swirling action creates a central recirculation zone (CRZ) and two external recirculation zones (ERZ) at the head of burner, which contributes to stabilizing combustion. Due to the wide variety of physical and chemical properties associated with IDF, previous research was often multi-faceted, focusing on key aspects, including burner design, flow-flame operation, flame appearance (such as size, color, and shape), and flame structure [3, 22, 23, 24]. Further, several studies [25, 26] have examined the stability and emission characteristics of dual flame structures, focusing on centerline temperature distribution and oxygen concentration in turbulent IDF. The combustion characteristics of IDF are heavily influenced by fuel/air mixing, making the fluid dynamics of the flame a crucial factor. In the laminar regime of IDF, fuel and oxidizers are transported to the reaction zone primarily through diffusion. In contrast, in turbulent regimes, the mixing of fuel and air occurs more vigorously due to fuel impingement and radial and tangential flows. As a result, IDF exhibits several dynamic characteristics. While many studies have been conducted on IDF combustors, their dynamics remain unexplored.
@@
Nalawade et al. [27] attempted to investigate the blowoff dynamics of turbulent IDF by synchronizing high speed CH* chemiluminescence imaging with pressure sensors. They found that blowoff happens due to higher strain rates caused by the simultaneous spatiotemporal interaction of turbulent flow structure with flame surface. Kabiraj et al.  [28] examined the nonlinear transitions from thermoacoustic oscillations to flame blowout due to intermittency, resulting from changes in the position of the combustion source within a ducted laminar premixed conical flame. They observed a subcritical Hopf bifurcation in the mean acoustic pressure oscillations as the combustor transitions into limit cycle oscillations. Additionally, the limit cycle oscillations displayed a bifurcation into quasi-periodic oscillations, leading to an intermittency state. Thus, the study  [28] explored the pathway to blowout via thermoacoustic instability in a single conical flame burner confined within a closed–open borosilicate glass duct.
@@
Further, Sen et al. [29] investigated the dynamics of a laboratory-scale ducted inverse diffusion flame with the variation of the position of flame within the duct. Their study revealed the presence of limit cycles, type-II intermittency, and homoclinic orbits through phase space reconstruction and recurrence analysis on the time series of acoustic signals. Recently, Bhattacharya et al. [30] investigated the onset of thermoacoustic instability using three recurrence network tools: global efficiency, average degree centrality, and global clustering coefficient. However, unlike the extensive research on premixed and non-premixed flames, transitions to thermoacoustic instability in IDF have received little attention which underscores the need for further research into the dynamics of IDF [12] with varying air and fuel flow rates.
@@
The dynamics associated with thermoacoustic instability are inherently non-linear, making it essential to investigate these transitions using non-linear methods [13, 18]. Illingworth et al. [31] tried to investigate the nonlinear thermoacoustic oscillations of a ducted Burke-Schumann diffusion flame by employing a combination of matrix-free continuation and flame-describing function (FDF) methods. Their study revealed that the inlet boundary conditions during the thermoacoustic cycle significantly contribute to the nonlinearity observed in the flame’s response to harmonic velocity fluctuations across a range of forcing frequencies and amplitudes. However, it is important to note that the flame describing function serves as an approximation, linearizing the nonlinear interactions between acoustic waves and flame dynamics. This simplification may neglect critical elements of complex flame behavior. On the other hand, the complexity of unsteady combustion behavior arises from the presence of multiple spatial and temporal scales. As a result, the fractal dimension of such systems can provide valuable insights into their complexity [32]. Typically, a single dimension can effectively describe the fractal dynamics of a system, leading to the concept of ”monofractal.” The idea of self-similarity, which is used to represent similar dynamics across different temporal and spatial scales in monofractals, was first introduced by Mandelbrot [33] in the study of turbulent flows. However, in complex systems, understanding their multiscale behavior often requires a range of fractal dimensions. This approach examines the spectrum of fractal dimensions and is known as multifractal analysis [34]. Mandelbrot first introduced the concept of multifractality in 1983 [35] and further developed it in 1989 [36]. Since then, the popularity of this concept has spread rapidly in various fields, including engineering [37], finance [38], geophysics [39], astrophysics [40], chemistry [41], and medical science [42, 43].
@@
Like different fields, in the combustion field, the multifractal approach has received significant attention. Niu et al. [44] confirmed the multifractal nature of the flame in a four-burner impinging entrained-flow gasifier by performing multifractal detrended fluctuation analysis on the high-speed flame images. Later, Sen et al. [45] investigated the dynamics of the cycle-to-cycle heat release rate variations in spark ignition engines at different spark advance angles and found a higher degree of multifractality of the dynamics at higher spark advance angles. In their subsequent study, Sen et al. [46] characterized the changes in the dynamics of a lean-fueled, multi-cylinder spark-ignition (SI) engine using multifractals and wavelets. Further, Curto-Risso et al. [47] used both monofractal and multifractal approaches to characterize the heat release rate fluctuations for different fuel-air ratios in a simulated spark ignition engine. They identified the presence of weak anticorrelations in the fluctuations for lean and stoichiometric conditions. Gotoda et al. [48] found the multifractal behavior of acoustic pressure oscillations in a swirl-stabilized turbulent combustor prior to the occurrence of a lean blowout. Nair and Sujith [49] noticed a scale invariance and multifractal behavior of unsteady pressure oscillations during the presence of combustion noise state in both swirl and bluff body stabilized combustors. They found that the multifractal behavior disappears at the onset of thermoacoustic instability in those combustors. Recently, Ragunathan et al. [32] studied the spatiotemporal dynamics of a bluff body stabilized combustor by performing multifractal analysis and discovered that multifractality exists in the flame topology for all dynamical phases during the transition to thermoacoustic instability. They found the fluctuations in multifractality spectrum at the onset of thermoacoustic instability that is nearly similar to the observation noticed by Sujith and Nair [49] in the temporal dynamics.
@@
Similar to these works mentioned in the last paragraph on conventional turbulent premixed combustion [48, 49, 32], the change in flow conditions or power input can induce thermoacoustic instability in IDF [29]. However, in contrast to conventional flames, the transitions leading to thermoacoustic instability in IDF caused by changes in control parameters are not well understood. Additionally, unlike the laminar and turbulent premixed flames [28], the development and testing of strategies for predicting the onset of thermoacoustic instability in IDF is limited.
@@
Hence, in light of the advancement in fractal theory, we apply multifractal detrended fluctuation analysis to examine the transitions to thermoacoustic instability in an IDF system under two different conditions: A. a constant power input (i.e., fuel flow rate) while Reynolds number is varied from laminar to turbulent. B. constant flow condition while power input is varied. We identify different regimes based on the variation in acoustic emissions captured by a microphone. Our analysis reveals a higher degree of multifractal behavior in the system dynamics at lower Reynolds numbers (with constant power input) and at lower power inputs (with constant Reynolds numbers). A loss of multifractality is captured when the system transitions to a higher Reynolds number (constant power input) and higher power input (constant Reynolds number). Further, we find that the Hurst exponent obtained through monofractal and multifractal spectrum tools, such as spectrum width and singular strength, are more effective than the conventional recurrence network matrices [30] in characterizing the dynamical regimes of the IDF system near thermoacoustic instability.
@@ III Methodology
The multifractal detrended fluctuation analysis (MFDFA) [34] can be implemented using the following steps: Let us consider a stochastic event, represented by a time series, x, with a length of N.
@@
MFDFA focuses on analyzing the characteristics of fluctuations in stochastic events. By subtracting the mean over whole time series [50], we create a modified time series ($x^{\prime}$) with a zero mean. This approach provides a clearer understanding of the profile of the time series.
@@
In general, MFDFA is employed to treat time series that are random walk in nature. This technique uncovers essential patterns, making it crucial for informed decision-making in data analysis. If the time series shows a noise-like structure instead of resembling a random walk, it must be transformed into a series that displays random walk characteristics before applying MFDFA. Integrating a noise-like series yields a random walk-like series that can then be treated with MFDFA for further analysis [51]. If $x^{\prime}$ is noise-like in nature, then, supposing $Z$ is the random walk derived from $x^{\prime}$:
@@
It should be noted that the first step of this process, i.e., the mean subtraction, is optional for a random walk-like series. However, it is compulsory to implement the same for noise-like series before performing the integration to convert it into a random walk. To distinguish between random-walk and noise-like time series, a potential solution proposed by Eke et al. [52] is to conduct a monofractal DFA prior to performing MFDFA. Time series exhibit noise-like characteristics when the Hurst exponent, $H$ (discussed later in the current section), falls between 0.2 and 0.8. In this scenario, MFDFA can be applied directly without the need for transformation of the time series. Conversely, when $H$ is within the range of 1.2 to 1.8, the time series behaves more like a random walk.
@@
where, $i$ = 2, 3, …, $m-1$. $S_{min}$ and $S_{max}$ are taken as $2^{9}$ and $2^{14}$. According to the thumb rule, $S_{max}$ is kept less $N$/10 to ensure at least 10 segments of $Z$ are formed, and MFDFA works accurately  [51].
@@
The length of the time series, $N$, in general, is not an integral multiple of the scales, $S_{i}$. As a result, a small segment of the series, $Z_{c}$, $Z_{c+1}$, …, $Z_{n}$, where $c=(S_{i}\times k_{i})+1$ and $m–k_{i}<S_{i}$ remains unexplored. In order to avoid ignoring this small segment, the same process is repeated starting from the end of the series or by reversing the series. As a result, for each scale, $S_{i}$, $2k_{i}$ number of segments are formed from the reversed series given by:
@@
Next, the $q$-orders are used to provide weightage for the local fluctuations in order to identify the self-similarity of the fluctuations at different time scales. The $q$-orders must have both positive and negative values that are not necessarily integers. The negative $q$-orders provide higher weightage to small-scale variations, while the positive $q$-orders render structure functions with higher weightage to large-scale variations. However, using extreme $q$-orders can lead to the dominance of fluctuations, which results in skewed and imprecise calculations. Therefore, it is important to be cautious when selecting the $q$-orders for a study. In most combustion studies, a $q$-order range of -5 to 5 has been commonly used [53].
@@
On a log-log plot, scales $S_{i}$ (where $i$ = 1, 2, …, $m$) are plotted on the abscissa and the structure function corresponding to each of these scales is plotted on the ordinate for the various $q$-orders considered. A linear best fit line is generated via least square method and the slope of this line gives us the generalized Hurst exponent($H(q)$) for various $q$. If $H(q)$ is dependent on $q$, the signal has multifractal characteristics. For positive values of $q$, $H(q)$ represents the scaling behavior of segments with large fluctuations, while it represents the scaling behavior of segments with small fluctuations for negative values of $q$. For a specific case where $q=2$, the term $H_{q}$ is commonly called the Hurst exponent [54].
@@
Using the generalized Hurst exponent, $H(q)$, the other MFDFA parameters are constructed. The scaling exponent or mass exponent, $\tau(q)$ is calculated as: $\tau(q)=qH(q)-1$. The singularity spectrum, $f(\alpha)$ is constructed from the scaling exponent as they are related via the Lengendre transform [55]:

### v3

@@ Abstract
Inverse diffusion flame, Thermoacosutic instability, Multifractal, Hurst exponent, Transition in dynamics.
preprint: AIP/123-QED
@@ I Introduction
“Rijke tube” Rijke (1859a, b) discovered by Professor P.L. Rijke in 1859, is a cylindrical tube with open ends that contains a heat source. This unique configuration generates two distinct types of flames, influenced by the introduction of oxidizer and fuel: normalized diffusion flame (NDF) Hunger et al. (2017) and inverse diffusion flame (IDF) De la Cruz-Ávila et al. (2017); Mikofski et al. (2006). In the NDF configuration, fuel is injected at the center, surrounded by an oxidizer. In contrast, the IDF configuration has a coaxial setup where the fuel surrounds a centralized flow of oxidizer. While both configurations provide stable operations, IDF is generally considered more suitable for practical applications because it results in lower emissions of thermal $NO_{x}$ and soot Sze et al. (2006); Choy et al. (2012); Kotb and Saad (2016, 2016), is less susceptible to lean blowout Sobiesiak and Wenzell (2005), and offers a wider stability margin Zhen et al. (2010). In IDF configuration, the central air jet stabilizes the flame and enhances flame stability by creating negative pressure along the flame axis.
@@
An IDF consists of two distinct zones: the flame torch and the flame base. These zones are connected by a section called the flame neck. In the flame neck region, intense mixing occurs due to the controlled velocity of the air jet, allowing air and unburned fuel to mix effectively before entering the flame torch zone Miao et al. (2016). Furthermore, an IDF combines the benefits of both normal diffusion flames and premixed flames. Therefore, the IDF concept can offer an alternative approach to both lean premixed combustion and normal diffusion flames in applications such as rocket engine combustion, staged combustion, industrial burners, and aviation Zhen et al. (2021). IDF is also beneficial in heating applications where a single flame needs to have a higher thermal load (i.e., a higher heat transfer rate) Zhen et al. (2021). Compared to typical diffusion flames or premixed flames, IDF is cleaner and safer for these types of applications.
@@
However, the IDF configuration converts heat into sound by promoting self-sustaining and self-amplifying standing waves. This phenomenon is highly relevant to the safety of combustion chambers in industrial burners Sujith and Unni (2020), rocket propulsion  De et al. (2024); Hashimoto et al. (2019); Kasthuri et al. (2022a), etc. The presence of high-amplitude acoustic waves from thermoacoustic instability can damage the structural integrity of engines Sujith and Unni (2020). This instability often manifests as loud noise, known as combustion roar, which is undesirable in both industrial and consumer applications and can lead to mission failures.  DeLaat and Paxson (2008). Additionally, thermoacoustic isntability can disrupt energy conversion efficiency, resulting in higher fuel consumption and lower output power. In some cases, it may also increase pollutant emissions, particularly in industrial combustion systems. Thus, the challenges of thermoacoustic instability in gas turbine engines, rocket motors, industrial boilers, oxygen-fuel Combustion and industrial process control remain significant for researchers in the twenty-first century Sujith and Pawar (2021); Silva (2023); Morgans and Yang (2024). Consequently, investigating thermoacoustic instability across various engine combustors, power generation systems, and aerospace applications has become a key area of focus for scientists. In this study, we examine the critical issues of thermoacoustic instability in an IDF combustor.
@@
Earlier, a significant amount of research had been conducted on IDF to understand their characteristics. Recently, Sun et al. Sun et al. (2024) employed the Reynolds-averaged Navier-Stokes (RANS) numerical method to analyze the combustion performance of an IDF combined with air swirling. Their results demonstrated that the swirling action creates a central recirculation zone (CRZ) and two external recirculation zones (ERZ) at the head of burner, which contributes to stabilizing combustion. Due to the wide variety of physical and chemical properties associated with IDF, previous research was often multi-faceted, focusing on key aspects, including burner design, flow-flame operation, flame appearance (such as size, color, and shape), and flame structure Hunger et al. (2017); Miao et al. (2014); Patel and Shah (2019); Zhu et al. (2018). Further, several studies Mahesh and Mishra (2008, 2010) have examined the stability and emission characteristics of dual flame structures, focusing on centerline temperature distribution and oxygen concentration in turbulent IDF. The combustion characteristics of IDF are heavily influenced by fuel/air mixing, making the fluid dynamics of the flame a crucial factor. In the laminar regime of IDF, fuel and oxidizers are transported to the reaction zone primarily through diffusion. In contrast, in turbulent regimes, the mixing of fuel and air occurs more vigorously due to fuel impingement and radial and tangential flows. As a result, IDF exhibits several dynamic characteristics. While many studies have been conducted on IDF combustors, their dynamics remain unexplored.
@@
Nalawade et al. Nalawade et al. (2023) attempted to investigate the blowoff dynamics of turbulent IDF by synchronizing high speed CH* chemiluminescence imaging with pressure sensors. They found that blowoff happens due to higher strain rates caused by the simultaneous spatiotemporal interaction of turbulent flow structure with flame surface. Kabiraj et al.  Kabiraj and Sujith (2012) examined the nonlinear transitions from thermoacoustic oscillations to flame blowout due to intermittency, resulting from changes in the position of the combustion source within a ducted laminar premixed conical flame. They observed a subcritical Hopf bifurcation in the mean acoustic pressure oscillations as the combustor transitions into limit cycle oscillations. Additionally, the limit cycle oscillations displayed a bifurcation into quasi-periodic oscillations, leading to an intermittency state. Thus, the study  Kabiraj and Sujith (2012) explored the pathway to blowout via thermoacoustic instability in a single conical flame burner confined within a closed–open borosilicate glass duct.
@@
Further, Sen et al. Sen et al. (2018) investigated the dynamics of a laboratory-scale ducted inverse diffusion flame with the variation of the position of flame within the duct. Their study revealed the presence of limit cycles, type-II intermittency, and homoclinic orbits through phase space reconstruction and recurrence analysis on the time series of acoustic signals. Recently, Bhattacharya et al. Bhattacharya et al. (2021) investigated the onset of thermoacoustic instability using three recurrence network tools: global efficiency, average degree centrality, and global clustering coefficient. However, unlike the extensive research on premixed and non-premixed flames, transitions to thermoacoustic instability in IDF have received little attention which underscores the need for further research into the dynamics of IDF Zhen et al. (2021) with varying air and fuel flow rates.
@@
The dynamics associated with thermoacoustic instability are inherently non-linear, making it essential to investigate these transitions using non-linear methods Sujith and Unni (2020); Sujith and Pawar (2021). Illingworth et al. Illingworth et al. (2013) tried to investigate the nonlinear thermoacoustic oscillations of a ducted Burke-Schumann diffusion flame by employing a combination of matrix-free continuation and flame-describing function (FDF) methods. Their study revealed that the inlet boundary conditions during the thermoacoustic cycle significantly contribute to the nonlinearity observed in the flame’s response to harmonic velocity fluctuations across a range of forcing frequencies and amplitudes. However, it is important to note that the flame describing function serves as an approximation, linearizing the nonlinear interactions between acoustic waves and flame dynamics. This simplification may neglect critical elements of complex flame behavior. On the other hand, the complexity of unsteady combustion behavior arises from the presence of multiple spatial and temporal scales. As a result, the fractal dimension of such systems can provide valuable insights into their complexity Raghunathan et al. (2020). Typically, a single dimension can effectively describe the fractal dynamics of a system, leading to the concept of "monofractal." The idea of self-similarity, which is used to represent similar dynamics across different temporal and spatial scales in monofractals, was first introduced by Mandelbrot Mandelbrot and Mandelbrot (1999) in the study of turbulent flows. However, in complex systems, understanding their multiscale behavior often requires a range of fractal dimensions. This approach examines the spectrum of fractal dimensions and is known as multifractal analysis Kantelhardt et al. (2002). Mandelbrot first introduced the concept of multifractality in 1983 Mandelbrot (1983) and further developed it in 1989 Mandelbrot (1989). Since then, the popularity of this concept has spread rapidly in various fields, including engineering Meneveau and Sreenivasan (1991), finance Morales et al. (2014), geophysics Telesca and Lapenna (2006), astrophysics Jones et al. (1988), chemistry Saha et al. (2020), and medical science Huynh et al. (2023); He et al. (2022).
@@
Like different fields, in the combustion field, the multifractal approach has received significant attention. Niu et al. Niu et al. (2008) confirmed the multifractal nature of the flame in a four-burner impinging entrained-flow gasifier by performing multifractal detrended fluctuation analysis on the high-speed flame images. Later, Sen et al. Sen et al. (2008) investigated the dynamics of the cycle-to-cycle heat release rate variations in spark ignition engines at different spark advance angles and found a higher degree of multifractality of the dynamics at higher spark advance angles. In their subsequent study, Sen et al. Sen et al. (2010) characterized the changes in the dynamics of a lean-fueled, multi-cylinder spark-ignition (SI) engine using multifractals and wavelets. Further, Curto-Risso et al. Curto-Risso et al. (2010) used both monofractal and multifractal approaches to characterize the heat release rate fluctuations for different fuel-air ratios in a simulated spark ignition engine. They identified the presence of weak anticorrelations in the fluctuations for lean and stoichiometric conditions. Gotoda et al. Gotoda et al. (2012) found the multifractal behavior of acoustic pressure oscillations in a swirl-stabilized turbulent combustor prior to the occurrence of a lean blowout. Nair and Sujith Nair and Sujith (2014) noticed a scale invariance and multifractal behavior of unsteady pressure oscillations during the presence of combustion noise state in both swirl and bluff body stabilized combustors. They found that the multifractal behavior disappears at the onset of thermoacoustic instability in those combustors. Recently, Ragunathan et al. Raghunathan et al. (2020) studied the spatiotemporal dynamics of a bluff body stabilized combustor by performing multifractal analysis and discovered that multifractality exists in the flame topology for all dynamical phases during the transition to thermoacoustic instability. They found the fluctuations in multifractality spectrum at the onset of thermoacoustic instability that is nearly similar to the observation noticed by Sujith and Nair Nair and Sujith (2014) in the temporal dynamics.
@@
Similar to these works mentioned in the last paragraph on conventional turbulent premixed combustion Gotoda et al. (2012); Nair and Sujith (2014); Raghunathan et al. (2020), the change in flow conditions or power input can induce thermoacoustic instability in IDF Sen et al. (2018). However, in contrast to conventional flames, the transitions leading to thermoacoustic instability in IDF caused by changes in control parameters are not well understood. Additionally, unlike the laminar and turbulent premixed flames Kabiraj and Sujith (2012), the development and testing of strategies for predicting the onset of thermoacoustic instability in IDF is limited.
@@
Hence, in light of the advancement in fractal theory, we apply multifractal detrended fluctuation analysis to examine the transitions to thermoacoustic instability in an IDF system under two different conditions: A. a constant power input (i.e., fuel flow rate) while Reynolds number is varied from laminar to turbulent. B. constant flow condition while power input is varied. We identify different regimes based on the variation in acoustic emissions captured by a microphone. Our analysis reveals a higher degree of multifractal behavior in the system dynamics at lower Reynolds numbers (with constant power input) and at lower power inputs (with constant Reynolds numbers). A loss of multifractality is captured when the system transitions to a higher Reynolds number (constant power input) and higher power input (constant Reynolds number). Further, we find that the Hurst exponent obtained through monofractal and multifractal spectrum tools, such as spectrum width and singular strength, are more effective than the conventional recurrence network matrices Bhattacharya et al. (2021) in characterizing the dynamical regimes of the IDF system near thermoacoustic instability.
@@ III Methodology
The multifractal detrended fluctuation analysis (MFDFA) Kantelhardt et al. (2002) can be implemented using the following steps: Let us consider a stochastic event, represented by a time series, x, with a length of N.
@@
MFDFA focuses on analyzing the characteristics of fluctuations in stochastic events. By subtracting the mean over whole time series Li et al. (2019), we create a modified time series ($x^{\prime}$) with a zero mean. This approach provides a clearer understanding of the profile of the time series.
@@
In general, MFDFA is employed to treat time series that are random walk in nature. This technique uncovers essential patterns, making it crucial for informed decision-making in data analysis. If the time series shows a noise-like structure instead of resembling a random walk, it must be transformed into a series that displays random walk characteristics before applying MFDFA. Integrating a noise-like series yields a random walk-like series that can then be treated with MFDFA for further analysis Ihlen (2012). If $x^{\prime}$ is noise-like in nature, then, supposing $Z$ is the random walk derived from $x^{\prime}$:
@@
It should be noted that the first step of this process, i.e., the mean subtraction, is optional for a random walk-like series. However, it is compulsory to implement the same for noise-like series before performing the integration to convert it into a random walk. To distinguish between random-walk and noise-like time series, a potential solution proposed by Eke et al. Eke et al. (2002) is to conduct a monofractal DFA prior to performing MFDFA. Time series exhibit noise-like characteristics when the Hurst exponent, $H$ (discussed later in the current section), falls between 0.2 and 0.8. In this scenario, MFDFA can be applied directly without the need for transformation of the time series. Conversely, when $H$ is within the range of 1.2 to 1.8, the time series behaves more like a random walk.
@@
where, $i$ = 2, 3, …, $m-1$. $S_{min}$ and $S_{max}$ are taken as $2^{9}$ and $2^{14}$. According to the thumb rule, $S_{max}$ is kept less $N$/10 to ensure at least 10 segments of $Z$ are formed, and MFDFA works accurately  Ihlen (2012).
@@
The length of the time series, $N$, in general, is not an integral multiple of the scales, $S_{i}$. As a result, a small segment of the series, $Z_{c}$, $Z_{c+1}$, …, $Z_{n}$, where $c=(S_{i}\times k_{i})+1$ and $m\textendash k_{i}<S_{i}$ remains unexplored. In order to avoid ignoring this small segment, the same process is repeated starting from the end of the series or by reversing the series. As a result, for each scale, $S_{i}$, $2k_{i}$ number of segments are formed from the reversed series given by:
@@
Next, the $q$-orders are used to provide weightage for the local fluctuations in order to identify the self-similarity of the fluctuations at different time scales. The $q$-orders must have both positive and negative values that are not necessarily integers. The negative $q$-orders provide higher weightage to small-scale variations, while the positive $q$-orders render structure functions with higher weightage to large-scale variations. However, using extreme $q$-orders can lead to the dominance of fluctuations, which results in skewed and imprecise calculations. Therefore, it is important to be cautious when selecting the $q$-orders for a study. In most combustion studies, a $q$-order range of -5 to 5 has been commonly used Kasthuri et al. (2019a).
@@
On a log-log plot, scales $S_{i}$ (where $i$ = 1, 2, …, $m$) are plotted on the abscissa and the structure function corresponding to each of these scales is plotted on the ordinate for the various $q$-orders considered. A linear best fit line is generated via least square method and the slope of this line gives us the generalized Hurst exponent($H(q)$) for various $q$. If $H(q)$ is dependent on $q$, the signal has multifractal characteristics. For positive values of $q$, $H(q)$ represents the scaling behavior of segments with large fluctuations, while it represents the scaling behavior of segments with small fluctuations for negative values of $q$. For a specific case where $q=2$, the term $H_{q}$ is commonly called the Hurst exponent Hurst (1951).
@@
Using the generalized Hurst exponent, $H(q)$, the other MFDFA parameters are constructed. The scaling exponent or mass exponent, $\tau(q)$ is calculated as: $\tau(q)=qH(q)-1$. The singularity spectrum, $f(\alpha)$ is constructed from the scaling exponent as they are related via the Lengendre transform Zia et al. (2009):